Advertisements
Advertisements
Question
Four fair coins are tossed once. Find the probability mass function, mean and variance for a number of heads that occurred
Advertisements
Solution
Let X be the random variable that denotes the number of heads when four coins are tossed once.
X ={0, 1, 2, 3}
n(S) = 16
| Values of random variable | 0 | 1 | 2 | 3 | 4 | Total |
| Number of elements in inverse image | 1 | 4 | 6 | 4 | 1 | 16 |
Probability mass function
| x | 0 | 1 | 2 | 3 | 4 |
| F(x) | `1/16` | `4/16` | `6/16` | `4/16` | `1/16` |
Mean: `mu = "E"("X")`
= `0 xx 1/16 + 1 xx 4/16 + 2 xx 6/6 + 3 xx 4/16 + 4 xx 1/16`
= `32/16`
= 2
Variance: `"E"("X"^2)`
= `0^2 xx 1/16 + 1^2 xx 4/16 + 2^2 xx 6/16 + 3^2 xx 4/16 + 4^2 xx 1/16`
= `0 + 4/16 + 24/16 + 36/16 + 16/16`
= `80/16`
= 5
Var(X) = E(X2) – [E(X)]2
= 5 – 4
= 1
APPEARS IN
RELATED QUESTIONS
For the random variable X with the given probability mass function as below, find the mean and variance.
`f(x) = {{:(2(x - 1), 1 < x ≤ 2),(0, "otherwise"):}`
If µ and σ2 are the mean and variance of the discrete random variable X and E(X + 3) = 10 and E(X + 3)2 = 116, find µ and σ2
The probability density function of the random variable X is given by
`f(x) = {{:(16x"e"^(-4x), x > 0),(0, x ≤ 0):}`
find the mean and variance of X
Choose the correct alternative:
Consider a game where the player tosses a six-sided fair die. If the face that comes up is 6, the player wins ₹ 36, otherwise he loses ₹ k2, where k is the face that comes up k = {1, 2, 3, 4, 5}. The expected amount to win at this game in ₹ is
Choose the correct alternative:
Four buses carrying 160 students from the same school arrive at a football stadium. The buses carry, respectively, 42, 36, 34, and 48 students. One of the students is randomly selected. Let X denote the number of students that were on the bus carrying the randomly selected student. One of the 4 bus drivers is also randomly selected. Let Y denote the number of students on that bus. Then E(X) and E(Y) respectively are
Choose the correct alternative:
If P(X = 0) = 1 – P(X = 1). If E[X] = 3 Var(X), then P(X = 0) is
Choose the correct alternative:
If X is a binomial random variable with I expected value 6 and variance 2.4, Then P(X = 5) is
What do you understand by Mathematical expectation?
How do you defi ne variance in terms of Mathematical expectation?
State the definition of Mathematical expectation using continuous random variable
A person tosses a coin and is to receive ₹ 4 for a head and is to pay ₹ 2 for a tail. Find the expectation and variance of his gains
Choose the correct alternative:
E[X – E(X)]2 is
Choose the correct alternative:
An expected value of a random variable is equal to it’s
Choose the correct alternative:
The distribution function F(x) is equal to
Prove that if E(X) = 0, then V(X) = E(X2)
What is the expected value of a game that works as follows: I flip a coin and if tails pay you ₹ 2; if heads pay you ₹ 1. In either case, I also pay you ₹ 0.50
The time to failure in thousands of hours of an important piece of electronic equipment used in a manufactured DVD player has the density function
f(x) = `{{:(2"e"^(-2x)",", x > 0),(0",", "otherwise"):}`
Find the expected life of this piece of equipment
